OpenStudy (anonymous):

I C dollars are deposited in an account paying r percent annual interest, the amount after x years is A=C(1=r)^x . Approximate the amount in the account after years, when C=250, r=23%, and x=21 years. A=$

OpenStudy (anonymous):

A=C(1=r)^x 1=r or 1+r?

OpenStudy (anonymous):

I think 1 +r

OpenStudy (anonymous):

A=250(1+0.23)^21 here the equation can you solve from now?

OpenStudy (anonymous):

A=250(1.23)^21 A=250*77.26936447 A=$19.317.34112 round $ 19.317